OpenStudy (anonymous):

Is the set of multiples of 4 closed under addition? Explain why or provide a counterexample if not. A. Yes, because the sum of any two multiples of 4 is also a multiple of 4. B. Yes, because the sum of any two multiples of 4 is also a multiple of 8. C. No, and a counterexample is –4 + (–8) = –12. D. No, and a counterexample is 26 + 16 = 42.

OpenStudy (anonymous):

@geerky42

OpenStudy (freckles):

well let's look at these "counterexamples" and see if they are actual counterexamples

OpenStudy (freckles):

are 4 and 8 multiples of 4?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

4 times two 8

OpenStudy (freckles):

is the the sum of -4+(-8) a multiple of 4?

OpenStudy (freckles):

that is is -12 a multiple of 4?

OpenStudy (anonymous):

umm...

OpenStudy (anonymous):

no?

OpenStudy (freckles):

4 times 3 is 12 so -12 is a multiple of 4

OpenStudy (anonymous):

oh XD im not that smart

OpenStudy (freckles):

so this is definitely not an example showing the multiples of 4 aren't closed under addition

OpenStudy (anonymous):

kk

OpenStudy (freckles):

so lets look at the other counterexample

OpenStudy (anonymous):

okay

OpenStudy (freckles):

is 26 a multiple of 4/

OpenStudy (anonymous):

no?

OpenStudy (freckles):

so we need to go no further there we definitely know that can not be a choice then

OpenStudy (freckles):

so the answer has to be yes

OpenStudy (freckles):

do you think it is A or B and why

OpenStudy (swissgirl):

4k+4k=8k=4(2k) where k is an integer

OpenStudy (anonymous):

whats an integer

OpenStudy (freckles):

integer={...,-3,-2,-1,0,1,2,3,....,100,...100000,....}

OpenStudy (anonymous):

dont get it.

OpenStudy (swissgirl):

an integer is a whole number so 1,2 3, or -1,-2 or -3 1/2 or 1.55 are not integers so fractions and decimals are NOT and integer

OpenStudy (anonymous):

oh ok

OpenStudy (freckles):

also I would probably write 4L+4K=4(L+K) where L+K is an integer

OpenStudy (freckles):

since K and L are integers

OpenStudy (swissgirl):

ohhh ya sorryyy . Forgot to take that into consideration

OpenStudy (swissgirl):

Ok so basically when you add two integers that are multiples of 4 the sum is a multiple of 4

OpenStudy (anonymous):

so a?

OpenStudy (swissgirl):

lol yaaa -.-

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

you're awesome @swissgirl

OpenStudy (swissgirl):

Not really sure abt that @freckles is the awesome one

OpenStudy (anonymous):

you're both awesome!